# Disturbance-Adaptive Data-Driven Predictive Control: Trading Comfort Violation for Savings in Building Climate Control

### v3

@@ I Introduction
The building sector accounts for approximately 34% of global energy consumption, with a significant portion attributed to Heating, Ventilation, and Air Conditioning (HVAC) systems [united2023]. These systems are typically managed by Building Management Systems (BMS) using classical control strategies, such as rule-based controllers [salsbury2005survey]. However, these traditional approaches often fall short in efficiency, presenting a clear opportunity to improve energy performance while maintaining occupant comfort.
@@
Model Predictive Control (MPC) has emerged as a promising approach to address this challenge [drgovna2020all, killian2016ten]. MPC leverages a predictive building model to optimize economic objectives while satisfying comfort constraints, with control actions applied in a receding-horizon manner. Compared to traditional methods, MPC demonstrates superior control performance and robustness in building climate control [balali2023energy, stoffel2023evaluation, pergantis2024field]. Nonetheless, implementing MPC faces many practical challenges, particularly in accurately modeling the complex dynamics of buildings [drgovna2020all, killian2016ten].
@@
To mitigate this modeling challenge, Data-Driven Predictive Control (DPC) using black-box or gray-box techniques has gained considerable attention [xiao2023building, stoffel2023evaluation, bunning2022physics]. Compared to MPC with physics-based models [jorissen2019taco], DPC requires less engineering effort and expert knowledge. For example, [lian2023adaptive] demonstrated that a linear DPC, developed using only two days of operational building data, achieved an 18% reduction in energy consumption compared to an industrial control method in a real-world application. Furthermore, [di2022lessons] highlighted the potential of different data-driven modeling methods for addressing specific building control scenarios through a pragmatic, qualitative comparison study. Despite these advances, both physics-based and data-driven approaches remain susceptible to model uncertainties and measurement noise, often resulting in suboptimal performance or comfort violations [drgovna2020all].
@@
To address uncertainties, robust MPC approaches have been proposed to ensure constraint satisfaction [bemporad2007robust]. Recent work combines robust MPC with data-driven methods to learn system dynamics and preserve robust constraint guarantees [alanwar2022robust, de2024koopman, dubied2025robust, schimperna2024robust]. For instance, [alanwar2022robust] proposes a data-driven robust MPC for unknown linear systems that constructs reachable sets directly from noisy input–output data, guaranteeing robust constraint satisfaction under bounded process and measurement noise. In building control, robust MPC has been integrated with black-box and gray-box models identified from data, including resistor-capacitor (RC) models [gao2023energy], Gaussian processes [maddalena2022experimental], and neural networks [li2022tube, mahmood2023data]. For example, [gao2023energy] identifies a gray-box RC model and designs a tube-based MPC that tightens constraints, reducing operating cost by at least 24% and improving the temperature tracking under multiple uncertain predictions. In a complementary direction, [hu2023multi] constructs a data-driven, clustered weather-forecast uncertainty set and embeds it in an affine-disturbance-feedback robust MPC, yielding up to 8.8% energy savings compared to conventional robust MPC approaches. However, their reliance on worst-case optimization can lead to excessive conservatism and higher costs.
@@
In contrast, stochastic MPC offers a more balanced approach by allowing occasional comfort violations [mesbah2016stochastic], utilizing chance constraints to meet comfort requirements with a predefined probability. For example, [oldewurtel2012use] demonstrated the application of applied stochastic MPC in various building cases, highlighting its capability to balance energy efficiency and occupant comfort through sensitivity analysis of probability parameters. Nevertheless, most stochastic MPC methods are still conservative due to sufficient conditional chance constraints[korda2012stochastic] and depend on precise knowledge of disturbance distributions, which is rarely feasible in real-world applications [mesbah2016stochastic]. This raises the critical question of how to effectively manage comfort violations to achieve cost savings without exact disturbance characterization.
@@
Recent efforts have explored stochastic MPC formulations that enforce average constraint violation bounds over time [korda2012stochastic, korda2014stochastic, fleming2017time, shi2024dad, oldewurtel2013adaptively, ghosh2025adaptive]. These methods adapt disturbance bounds or constraint tightening based on closed-loop violations, directly related to the comfort criteria in building climate control. The authors of [korda2012stochastic] demonstrated that such approaches could reduce control costs compared to traditional stochastic MPC. More recently, [shi2024dad] proposed a disturbance-adaptive scheme that ensures asymptotic and robust average violation bounds, even with inaccurate disturbance quantification. However, applying existing average violation guarantees to building climate control is challenging because the guarantees rely on assumptions that are difficult to satisfy or implement in buildings. In particular, several results ([korda2012stochastic, korda2014stochastic, fleming2017time, shi2024dad]) assume the existence of a robust invariant set computed from a state-space model. Because buildings are modeled in various ways (such as adaptive data-driven predictors) and are driven by time-varying exogenous inputs (such as weather and occupancy), computing such sets without excessive conservatism can be challenging. Other results ([oldewurtel2013adaptively, ghosh2025adaptive]) assume a control policy with a one-step corrective property: if the average violation so far is too large or too small, the next-step violation will certainly disappear or occur. This is a strong assumption under actuator saturation, duty-cycling, and the large thermal inertia and slow thermal dynamics of buildings. Furthermore, while some studies have explored building control applications [oldewurtel2013adaptively, korda2014stochastic], these have primarily used simple linear time-invariant models, lacking validation in high-fidelity simulations or real-world experiments.
@@
This paper extends the disturbance adaptive method [shi2024dad] to a DPC framework without computing a robust invariant set, enabling a trade-off between the average comfort violations and cost savings in its application of building climate control. The key contributions are as follows:
@@ II-B Problem Statement
This paper focuses on building climate control, aiming to enhance the energy efficiency of HVAC systems while maintaining comfortable indoor temperature levels [oldewurtel2012use].
@@
These constraints relax the temperature range during non-working hours to save energy. Notably, occasional violations of comfort constraints are permitted under certain building regulations [comite2007indoor, sourbron2011evaluation]. To account for this, and as a central focus of this paper, we formalize an asymptotic average comfort violation bound as follows:
@@
In the constraint (2), $\alpha\in(0,1]$ specifies the allowable average violation bound. According to [sourbron2011evaluation], the value of $\alpha$ should be determined through pre-design discussions between building designers and clients. As shown in [oldewurtel2012use], increasing $\alpha$ results in higher comfort violations for greater energy savings, which was validated using a stochastic MPC method. However, the approach relied on sensitivity analysis of a probability parameter to satisfy the desired $\alpha$-level bound for average violations in the closed loop [oldewurtel2012use]. This limitation is common in stochastic MPC methods, where fixed conditional chance constraints introduce excessive conservatism and practical estimates of disturbance distributions are often quite inaccurate.
@@ III-A Disturbance-Adaptive Data-Driven Predictive Control
(i) One option is to design $\pi^{B}$ to practically stabilize the system to a constraint-admissible set. Choose a compact subset of its domain of attraction, and define $\mathcal{Y}_{\text{lim}}$ as a robust backward reachable set of this subset. For linear systems, [shi2024dad] and [korda2012stochastic] present a systematic approach based on robust control invariant sets. This design ensures that whenever $\pi^{B}$ is activated, constraint violations vanish after a user-defined finite number of steps, thereby satisfying Property 1.
@@ III-B Guarantees by DAD-DPC
First, we establish that at time $t$, the closed-loop average violation under the DAD-MPC is bounded related to the minimal and maximal $\alpha_{t}$ from initial time, denoted as $\alpha_{min,t}:=\min_{i\in\operatorname{\mathbb{N}}_{0}^{t}}\alpha_{i},\alpha_{max,t}:=\max_{i\in\operatorname{\mathbb{N}}_{0}^{t}}\alpha_{i}$, respectively [shi2024dad].
@@ III-D1 Data-driven Predictive Control
[willems2005note, Theorem 1] Consider a controllable linear system, $x_{t+1}=Ax_{t}+Bu_{t}$ and $y_{t}=Cx_{t}+Du_{t}$, with $n_{x}$ as the system’s order. If $\{u_{i}\}_{i=1}^{T}$ is persistently exciting of order $L+n_{x}$, then $\text{colspan}(\begin{bmatrix}H_{u}^{\top}&H_{y}^{\top}\end{bmatrix}^{\top})=\mathfrak{B}_{L}(A,B,C,D)$, where $\mathfrak{B}_{L}(A,B,C,D)$ represents the set of all $L$-step trajectories produced by the linear system.
Although this lemma is originally established for noise-free linear systems, extensions to systems with noisy measurements and nonlinear dynamics have been developed [dorfler2022bridging, lian2023adaptive, yin2024data]. Furthermore, many extensions have been experimentally validated in various building types for diverse climate control tasks [chinde2022data, lian2023adaptive, shi2025adaptive, yin2024data].
@@
In the low-level optimization problem, the decision variables are $g_{l}\in\mathbb{R}^{T-L+1},\delta_{y}\in\mathbb{R}^{t_{init}n_{y}}$, where $n_{y}$ is the output dimension. The regularization cost with the parameter $Q_{g}$ is used to mitigate the effects of measurement noise and system nonlinearity [lian2023adaptive]. Given the previous measurements $\mathbf{u}_{init},\;\mathbf{y}_{init}$ and $\mathbf{w}_{init}$, the equations (8c) and (8d) predict the output sequence $\mathbf{y}_{pred}$ for specified control inputs $\mathbf{u}_{pred}$ and predictive external inputs $\mathbf{w}_{pred}$.
In the high-level optimization problem, the controller optimizes the user-defined control objective subject to input and output constraints. This RB-DPC directly applies constraint tightening on the output prediction $\mathbf{y}_{pred}$, as shown in (8a) (similar to [berberich2020robust]). The extent of constraint tightening is determined by the disturbance bound estimator $\mathcal{D}_{i}(\sigma)$ and the confidence level $\sigma$. In this paper, we design $\mathcal{D}_{i}(\sigma)$ using conformal prediction, a method detailed in the next section. This approach enables a data-driven estimation of overall output uncertainties arising from various sources, including model uncertainties and measurement noise. In contrast, previous works such as [lian2023adaptive, shi2025adaptive] primarily address robustness against weather forecast errors. In addition, to ensure recursive feasibility, soft output constraints are included, with a quadratic penalty applied using the weight $Q_{\delta}\in\operatorname{\mathbb{R}}$. This parameter also allows the controller to adjust violations, particularly in the nominal case $\pi(z_{t},\mathcal{D}(1))$.
This bi-level DPC formulation (8) is well-suited for building climate control due to several advantages. It is insensitive to the choice of model order and does not require an observer [yin2024data]. Its linear structure ensures both data and computational efficiency. Compared to standard single-level DPC formulations based on Willems’ Fundamental Lemma, the bi-level framework significantly reduces tuning efforts [shi2025adaptive]. Moreover, it supports efficient adaptive updates, allowing it to approximate varying building operational conditions and enhance representational capability [shi2023efficient]. Empirical validations further demonstrate the effectiveness of this bi-level DPC [lian2023adaptive, shi2025adaptive]. For example, [shi2025adaptive] employed a similar DPC scheme using ten days of operational data for a two-month demand response task, achieving a 29% cost compared to an industrial rule-based method. Therefore, these attributes highlight the bi-level DPC as a practical and efficient solution for building climate control with low commissioning efforts.
@@ III-D2 Disturbance quantification using Conformal Prediction
Conformal prediction is a distribution-free, data-driven method for generating prediction intervals for regression models [cp_vovk2005, cp_intro_angelopoulos2021]. In this paper, we employ Split Conformal Prediction (SCP[cp_vovk2005, cp_intro_angelopoulos2021] to construct $\mathcal{D}_{i}(\sigma)$.
@@ III-D3 The resulting DAD-DPC
The disturbance bound estimator (11) derived from conformal prediction introduces box constraints on $d_{i}$ in the RB-DPC (8). When polyhedral output constraints, such as the comfort constraint (1) commonly used in buildings climate control, are applied, the computational complexity of $\pi(z_{t},\mathcal{D}(\sigma))$ remains equivalent to that of the nominal controller $\pi(z_{t},\mathcal{D}(1))$. If, in addition, the input constraint set $\mathcal{U}$ is polyhedral and $I(y_{i},u_{i})$ is a positive-definite quadratic function or an affine function, (8) is quadratic programming (QP), which can be efficiently solved using QP solvers such as Gurobi [gurobi] and MATLAB Quadprog [matlab:quadprog].
@@ IV-A Setup of simulation cases
The simulation study leverages the high-fidelity Building Optimization Testing Framework (BOPTEST) [blum2021building], which provides diverse cases for benchmarking building control performance. Three cases with distinct configurations were deployed to validate the DAD-DPC method. The setups of these cases, including building architecture, HVAC systems, occupancy schedules, and climate data, are summarized in Table I. Further details for these emulators are publicly available. Case 1 (“BESTEST hydronic heat pump”) is provided directly by BOPTEST [^1] https://ibpsa.github.io/project1-boptest/testcases/index.html. Case 2 (“single zone commercial hydronic”) and Case 3 (“Multi-zone office floor”) were provided through the BOPTEST Challenge [boptestChallenge] [^2] https://adrenalin.energy/BOPTEST-Challenge-Smart-building-HVAC-control. Specifically, Case 2 is a modified version of the original “single zone commercial hydronic,” and Case 3 is a modified version of the original “multizone office simple air”11footnotemark: 1. Both modified FMUs and their descriptions are available online [^3] https://data.sintef.no/feature/fe-7d9861ec-ae6e-4947-ae43-85f27fee3cd3 and can be imported into the BOPTEST package as described in the Challenge instructions.
@@
Figure 3: Illustration of five zones in Case 3 [deru2008doe].

### v4

@@ I Introduction
The building sector accounts for approximately 34% of global energy consumption, with a significant portion attributed to Heating, Ventilation, and Air Conditioning (HVAC) systems [38]. These systems are typically managed by Building Management Systems (BMS) using classical control strategies, such as rule-based controllers [39]. However, these traditional approaches often fall short in efficiency, presenting a clear opportunity to improve energy performance while maintaining occupant comfort.
@@
Model Predictive Control (MPC) has emerged as a promising approach to address this challenge [15, 25]. MPC leverages a predictive building model to optimize economic objectives while satisfying comfort constraints, with control actions applied in a receding-horizon manner. Compared to traditional methods, MPC demonstrates superior control performance and robustness in building climate control [3, 45, 37]. Nonetheless, implementing MPC faces many practical challenges, particularly in accurately modeling the complex dynamics of buildings [15, 25].
@@
To mitigate this modeling challenge, Data-Driven Predictive Control (DPC) using data-driven or gray-box techniques has gained considerable attention [50, 45, 9]. Compared to MPC with physics-based models [24], DPC requires less engineering effort and expert knowledge. For example, [29] demonstrated that a linear DPC, developed using only two days of operational building data, achieved an 18% reduction in energy consumption compared to an industrial control method in a real-world application. Furthermore, [13] highlighted the potential of different data-driven modeling methods for addressing specific building control scenarios through a pragmatic, qualitative comparison study. Despite these advances, both physics-based and data-driven approaches remain susceptible to model uncertainties and measurement noise, often resulting in suboptimal performance or comfort violations [15].
@@
To address uncertainties, robust MPC approaches have been proposed to ensure constraint satisfaction [4]. Recent work combines robust MPC with data-driven methods to learn system dynamics and preserve robust constraint guarantees [1, 12, 16, 40]. For instance, [1] proposes a data-driven robust MPC for unknown linear systems that constructs reachable sets directly from noisy input–output data, guaranteeing robust constraint satisfaction under bounded process and measurement noise. In building control, robust MPC has been integrated with data-driven and gray-box models identified from data, including resistor-capacitor (RC) models [19], Gaussian processes [30], and neural networks [28, 31]. For example, [19] identifies a gray-box RC model and designs a tube-based MPC that tightens constraints, reducing operating cost by at least 24% and improving the temperature tracking under multiple uncertain predictions. In a complementary direction, [22] constructs a data-driven, clustered weather-forecast uncertainty set and embeds it in an affine-disturbance-feedback robust MPC, yielding up to 8.8% energy savings compared to conventional robust MPC approaches. However, their reliance on worst-case optimization can lead to excessive conservatism and higher costs.
@@
In contrast, stochastic MPC offers a more balanced approach by allowing occasional comfort violations [33], utilizing chance constraints to meet comfort requirements with a predefined probability. For example, [35] demonstrated the application of applied stochastic MPC in various building cases, highlighting its capability to balance energy efficiency and occupant comfort through sensitivity analysis of probability parameters. Nevertheless, most stochastic MPC methods are still conservative due to sufficient conditional chance constraints[26] and depend on precise knowledge of disturbance distributions, which is rarely feasible in real-world applications [33]. This raises the critical question of how to effectively manage comfort violations to achieve cost savings without exact disturbance characterization.
@@
Recent efforts have explored stochastic MPC formulations that enforce average constraint violation bounds over time [26, 27, 18, 41, 36, 20]. These methods adapt disturbance bounds or constraint tightening based on closed-loop violations, directly related to the comfort criteria in building climate control. The authors of [26] demonstrated that such approaches could reduce control costs compared to traditional stochastic MPC. More recently, [41] proposed a disturbance-adaptive scheme that ensures asymptotic and robust average violation bounds, even with inaccurate disturbance quantification. However, applying existing average violation guarantees to building climate control is challenging because the guarantees rely on assumptions that are difficult to satisfy or implement in buildings. In particular, several results ([26, 27, 18, 41]) assume the existence of a robust invariant set computed from a state-space model. Because buildings are modeled in various ways (such as adaptive data-driven predictors) and are driven by time-varying exogenous inputs (such as weather and occupancy), computing such sets without excessive conservatism can be challenging. Other results ([36, 20]) assume a control policy with a one-step corrective property: if the average violation so far is too large or too small, the next-step violation will certainly disappear or occur. This is a strong assumption under actuator saturation, duty-cycling, and the large thermal inertia and slow thermal dynamics of buildings. Furthermore, while some studies have explored building control applications [36, 27], these have primarily used simple linear time-invariant models, lacking validation in high-fidelity simulations or real-world experiments.
@@
This paper extends the disturbance adaptive method [41] to a DPC framework without computing a robust invariant set, enabling a trade-off between the average comfort violations and cost savings in its application of building climate control. The key contributions are as follows:
@@ II-B Problem Statement
This paper focuses on building climate control, aiming to enhance the energy efficiency of HVAC systems while maintaining comfortable indoor temperature levels [35].
@@
These constraints relax the temperature range during non-working hours to save energy. Notably, occasional violations of comfort constraints are permitted under certain building regulations [11, 44]. To account for this, and as a central focus of this paper, we formalize an asymptotic average comfort violation bound as follows:
@@
In the constraint (2), $\alpha\in(0,1]$ specifies the allowable average violation bound. According to [44], the value of $\alpha$ should be determined through pre-design discussions between building designers and clients. As shown in [35], increasing $\alpha$ results in higher comfort violations for greater energy savings, which was validated using a stochastic MPC method. However, the approach relied on sensitivity analysis of a probability parameter to satisfy the desired $\alpha$-level bound for average violations in the closed loop [35]. This limitation is common in stochastic MPC methods, where fixed conditional chance constraints introduce excessive conservatism and practical estimates of disturbance distributions are often quite inaccurate.
@@ III-A Disturbance-Adaptive Data-Driven Predictive Control
(i) One option is to design $\pi^{B}$ to practically stabilize the system to a constraint-admissible set. Choose a compact subset of its domain of attraction, and define $\mathcal{Y}_{\text{lim}}$ as a robust backward reachable set of this subset. For linear systems, [41] and [26] present a systematic approach based on robust control invariant sets. This design ensures that whenever $\pi^{B}$ is activated, constraint violations vanish after a user-defined finite number of steps, thereby satisfying Property 1.
@@ III-B Guarantees by DAD-DPC
First, we establish that at time $t$, the closed-loop average violation under the DAD-MPC is bounded related to the minimal and maximal $\alpha_{t}$ from initial time, denoted as $\alpha_{min,t}:=\min_{i\in\operatorname{\mathbb{N}}_{0}^{t}}\alpha_{i},\alpha_{max,t}:=\max_{i\in\operatorname{\mathbb{N}}_{0}^{t}}\alpha_{i}$, respectively [41].
@@ III-D1 Data-driven Predictive Control
[49, Theorem 1] Consider a controllable linear system, $x_{t+1}=Ax_{t}+Bu_{t}$ and $y_{t}=Cx_{t}+Du_{t}$, with $n_{x}$ as the system’s order. If $\{u_{i}\}_{i=1}^{T}$ is persistently exciting of order $L+n_{x}$, then $\text{colspan}(\begin{bmatrix}H_{u}^{\top}&H_{y}^{\top}\end{bmatrix}^{\top})=\mathfrak{B}_{L}(A,B,C,D)$, where $\mathfrak{B}_{L}(A,B,C,D)$ represents the set of all $L$-step trajectories produced by the linear system.
Although this lemma is originally established for noise-free linear systems, extensions to systems with noisy measurements and nonlinear dynamics have been developed [14, 29, 51]. Furthermore, many extensions have been experimentally validated in various building types for diverse climate control tasks [10, 29, 43, 51].
@@
In the low-level optimization problem, the decision variables are $g_{l}\in\mathbb{R}^{T-L+1},\delta_{y}\in\mathbb{R}^{t_{init}n_{y}}$, where $n_{y}$ is the output dimension. The regularization cost with the parameter $Q_{g}$ is used to mitigate the effects of measurement noise and system nonlinearity [29]. Given the previous measurements $\mathbf{u}_{init},\;\mathbf{y}_{init}$ and $\mathbf{w}_{init}$, the equations (8c) and (8d) predict the output sequence $\mathbf{y}_{pred}$ for specified control inputs $\mathbf{u}_{pred}$ and predictive external inputs $\mathbf{w}_{pred}$.
In the high-level optimization problem, the controller optimizes the user-defined control objective subject to input and output constraints. This RB-DPC directly applies constraint tightening on the output prediction $\mathbf{y}_{pred}$, as shown in (8a) (similar to [5]). The extent of constraint tightening is determined by the disturbance bound estimator $\mathcal{D}_{i}(\sigma)$ and the confidence level $\sigma$. In this paper, we design $\mathcal{D}_{i}(\sigma)$ using conformal prediction, a method detailed in the next section. This approach enables a data-driven estimation of overall output uncertainties arising from various sources, including model uncertainties and measurement noise. In contrast, previous works such as [29, 43] primarily address robustness against weather forecast errors. In addition, to ensure recursive feasibility, soft output constraints are included, with a quadratic penalty applied using the weight $Q_{\delta}\in\operatorname{\mathbb{R}}$. This parameter also allows the controller to adjust violations, particularly in the nominal case $\pi(z_{t},\mathcal{D}(1))$.
This bi-level DPC formulation (8) is well-suited for building climate control due to several advantages. It is insensitive to the choice of model order and does not require an observer [51]. Its linear structure ensures both data and computational efficiency. Compared to standard single-level DPC formulations based on Willems’ Fundamental Lemma, the bi-level framework significantly reduces tuning efforts [43]. Moreover, it supports efficient adaptive updates, allowing it to approximate varying building operational conditions and enhance representational capability [42]. Empirical validations further demonstrate the effectiveness of this bi-level DPC [29, 43]. For example, [43] employed a similar DPC scheme using ten days of operational data for a two-month demand response task, achieving a 29% cost compared to an industrial rule-based method. Therefore, these attributes highlight the bi-level DPC as a practical and efficient solution for building climate control with low commissioning efforts.
@@ III-D2 Disturbance quantification using Conformal Prediction
Conformal prediction is a distribution-free, data-driven method for generating prediction intervals for regression models [48, 2]. In this paper, we employ Split Conformal Prediction (SCP[48, 2] to construct $\mathcal{D}_{i}(\sigma)$.
@@ III-D3 The resulting DAD-DPC
The disturbance bound estimator (11) derived from conformal prediction introduces box constraints on $d_{i}$ in the RB-DPC (8). When polyhedral output constraints, such as the comfort constraint (1) commonly used in buildings climate control, are applied, the computational complexity of $\pi(z_{t},\mathcal{D}(\sigma))$ remains equivalent to that of the nominal controller $\pi(z_{t},\mathcal{D}(1))$. If, in addition, the input constraint set $\mathcal{U}$ is polyhedral and $I(y_{i},u_{i})$ is a positive-definite quadratic function or an affine function, (8) is quadratic programming (QP), which can be efficiently solved using QP solvers such as Gurobi [21] and MATLAB Quadprog [32].
@@ IV-A Setup of simulation cases
The simulation study leverages the high-fidelity Building Optimization Testing Framework (BOPTEST) [6], which provides diverse cases for benchmarking building control performance. Three cases with distinct configurations were deployed to validate the DAD-DPC method. The setups of these cases, including building architecture, HVAC systems, occupancy schedules, and climate data, are summarized in Table I. Further details for these emulators are publicly available. Case 1 (“BESTEST hydronic heat pump”) is provided directly by BOPTEST [^1] https://ibpsa.github.io/project1-boptest/testcases/index.html. Case 2 (“single zone commercial hydronic”) and Case 3 (“Multi-zone office floor”) were provided through the BOPTEST Challenge [7] [^2] https://adrenalin.energy/BOPTEST-Challenge-Smart-building-HVAC-control. Specifically, Case 2 is a modified version of the original “single zone commercial hydronic,” and Case 3 is a modified version of the original “multizone office simple air”11footnotemark: 1. Both modified FMUs and their descriptions are available online [^3] https://data.sintef.no/feature/fe-7d9861ec-ae6e-4947-ae43-85f27fee3cd3 and can be imported into the BOPTEST package as described in the Challenge instructions.
@@
Figure 3: Illustration of five zones in Case 3 [8].